OpenStudy (anonymous):

Quick!!! ....how to factor the sum or difference of cubes in a pynomial??... 2x^3+216

OpenStudy (anonymous):

.... is that your final answer

OpenStudy (anonymous):

a^3 - b^3 = (a - b)(a^2 + ab + b^2)

OpenStudy (anonymous):

a^3 + b^3 = (a + b)(a^2 - ab + b^2)

OpenStudy (anonymous):

These are the identities to be used...

OpenStudy (anonymous):

2x^3+216 \[(\sqrt{2}x)^{3} + 6^{3} = (\sqrt{2}x + 6)[(\sqrt{2}x)^{2} - (\sqrt{2}x)(6) + 6^{2}]\]

OpenStudy (anonymous):

= \[(\sqrt{2}x + 6)(2x ^{2} - 6\sqrt{2}x + 36)\]